# Machine learning applications in archaeological practices: a review

### v3

@@ 1 Introduction
Unlike traditional statistical models, like linear models such as ordinary least squares or general linear models that find solutions with formal equations, machine learning methods seek to use the data itself to create a model able to accurately evaluate new, unseen data. Machine learning is defined by Alpaydin as “programming computers to optimize a performance criterion using example data or past experience(Alpaydin, 2014, p.3). Machine learning models require a training phase, where the computer learns from the data to improve its model and thus its predictions. Once trained, machine learning models often produce faster and more accurate results than traditional methods with lower cost, often whilst also producing more accurate predictions [e.g. Kochkov et al. 2024]. This benefit can make machine learning an appealing solution to certain scientific questions (Verhagen, 2007; Hansen and Nebel, 2020; Calder et al., 2022; Brandsen and Koole, 2022; Orellana Figueroa et al., ), Orellana Figueroa et al., in press. However, this comes at the cost of both the models being difficult to interpret (Carvalho et al., 2019), and the training process generally requiring large amounts of (high quality) data, time, and compute power; in a way, front-loading the time and costs to the training process to produce fast and accurate predictions once trained (Sevilla et al., 2022).
@@
The frequency, extent, and success of machine learning applications across archaeology remain largely unknown and their implications and ethical considerations are under debate (Davis, 2020b; Bickler, 2021; Tenzer et al., 2024; Gattiglia, 2025; Neri and Dadà, 2025). Previous reviews of machine learning applications in archaeology mainly focused on remote sensing (Jamil et al., 2022; Argyrou and Agapiou, 2022), text analysis (Gonzalez-Perez et al., 2023), classification of ceramics (Ling et al., 2024), wear-use analysis (Eleftheriadou et al., 2025) and other artefacts (Naso and Sciuto, 2025), or only on a small set of study cases (Mantovan and Nanni, 2020). On the other hand, while Cacciari and Pocobelli [2022] have assessed machine learning applications across a wide range of fields of archaeological research, they mainly provide qualitative observations rather than a quantitative analysis as presented here, as is also the case for the overview published by Calder et al. [2022]. Palacios [2023] provided an innovative introduction on the status of machine learning in archaeology, although mainly focused on Bayesian approaches, and with a limited selection of reviewed papers mainly on archaeological predictive models. A recent comparison of machine learning methods to PCA and LDA for species identity from faunal material find that methods like RF and LDA outperform PCA (Cole et al., 2022).
@@ 2.1 Search strategy
The search strategy comprised two different and independent protocols: one performed for the automatic screening protocol, consisting of multiple searches, and one performed for the manual screening protocol, consisting of a single search [Table 1, Fig. 1]. The searches for the automatic protocol were performed by M.B and consisted of twelve queries each containing a combination of keywords [Box 1]. These keywords, while not covering the entire scope of possible archaeological study cases using machine learning (such as those not using the exact term of "machine learning"), already grasp a wide overview of machine learning applications in the field of archaeology. It is however important to note that this might have biased the search in favour of articles that use the buzzwords of “machine learning”. On the other hand, an earlier search using the manual protocol involved multiple queries using the 10 machine learning method families we used for our data analysis [see supplementary file], but these returned a large number of duplicate results, most of which were already captured by simply using "machine learning". The searches were performed in five online portals. All searches were conducted in English, except for the German National Library, where the search was conducted in German with only six of the queries [Box 1]. These searches covered all records published before 1 January 2023. In total, 1460 records were retrieved from the automated search protocol, of which 558 were unique.
@@ 2.2 Screening
Following the automatic and manual protocol searches, two independent screening procedures were performed on the retrieved records [Fig. 1]. For the automatic protocol, after excluding publications not written in English and records not published in scientific journals, the screening was performed with a script written in R 4.4.1 (R Core team, 2024) with the corpus and tm packages (Huang et al., 2021; Feinerer and Hornik, 2023) divided in two filters. The first screening step retrieved all documents whose full text contained one of the archaeology or machine learning keywords we defined for this filtering step [supplementary file]. The second screening step was performed in the same manner with the same keywords, but using only the abstract and title of the results of the first screening step that were found to contain the keywords in the full text. A total of 135 records were included from the automatic search and screening protocol.
The manual screening protocol consisted of additional steps. First, a preliminary exclusion was performed, which checked whether the publications were inaccessible, whether they were preprints that were not yet in print, and whether the record was for an entire book, amongst other criteria [Fig. 1,supplementary file]. Further screening was then performed based on the exclusion and inclusion criteria [supplementary file], firstly by examining the publication title alone. If after this, an article could not be securely excluded or included, a second screening step followed, now examining the abstract as well. If this was yet still insufficient for a secure inclusion or exclusion, the full-text was examined. A total of 93 records were included from the manual protocol search and screening protocol.
Both screening strategies used the same inclusion and exclusion criteria [Fig. 1,supplementary file], though the automatic screening script, by necessity, used only a simplified version that only involved the use of keywords, while the manual protocol search was done through the human-based analysis of text.
@@ 2.3 Data extraction
We systematically divided each of the included articles into one or more study cases if different applications (e.g. using different data or seeking different goals) were attested in the publication. Therefore, we obtained a total number of study cases (n = 147) superior to the sum of all included articles (n = 135). From all these study cases, we systematically recorded a total of nine features composed of different categories [Table 2]. These features were based on the authors’ ad hoc evaluation of possible features of interest in the evaluation of the study cases. No inter-rater reliability was calculated for the selection of these features. However, only four of these were needed to gain a clear overview of the practice of machine learning in archaeology: the architecture of the machine learning model used as well as its evaluation process (see below), the archaeological subfields where these methods were applied, and the type of task it was applied to [Fig. 2]. Though important, the other groups of features did not allow us to identify common patterns in the application of machine learning in archaeology. Thus, we included them only sporadically in our discussion. Furthermore, we extracted metadata related to the publication for our quantitative analysis. These metadata included the year of publication, the list of authors, the country of affiliation of the first author, the name of the journal, and whether the article was published through an open access modality.
Based on the model architecture, we divided machine learning methods into nine broad categories based on the family of algorithms they belonged to [Eleftheriadou et al. 2025, Fig. 2, supplementary file]. We granularly recorded the different types of models used, which were then classified into the relevant architecture category, but we also recorded the frequency of their use. Each model was classified into only one architecture category, even in cases where multiple categories may have been warranted. For example, random forest is an ensemble of decision trees, fitting into both “ensemble learning” and “decision trees”, but was classified into “ensemble learning” only, as that was the more relevant aspect of the algorithm.
Furthermore, according to the goals of the model application, we grouped the evaluation into three broad categories: classification, regression, and clustering [Fig. 2; Alpaydin 2014, pp. 5-13].
@@
Due to the wide range of subjects, we divided the field of archaeology into fifteen subfields [Fig. 2] based on Kelly and Thomas [2017]. The assignment of a study case to a specific subfield was based on the authors’ own evaluation [supplementary file]. No inter-rater reliability calculation was performed for the assignment. A single publication could be classified into several subfields based on their research questions or objectives, and therefore, the sum total entries in the subfield feature does not equal the total number of study cases included in the review.
Finally, we classified every article’s application into nineteen a posteriori categories based on the broader task (e.g. automatic structure detection) for which the machine learning methods were applied [Fig. 2]. These task categories were ad hoc and based on the authors’ evaluation of the corpus. Initially, we sought to define the tasks for which machine learning was applied for each publication from a very granular level, then aggregating the more granular task categories into broader ones a total of five times until we arrived at the nineteen categories reported here [supplementary file]. No inter-rater reliability calculation was performed for the creation of these tasks categories. These task categories were one of the most important characteristics to analyse, as these classes helped us understand not only the goals of the authors, but also the possible models that were more well-suited for the desired outcomes, based especially on previous applications.
@@ 3.1 Screening results
Since the two screening protocols were performed independently, we will report them as such, even if the records obtained after both screenings were merged into a single set of publications [Fig. 1]. Out of the 558 unique records obtained from the automatic protocol, four could not be screened with our script (0.72%, reason 1), another four were excluded, as they were not written in English (0.72%, reason 2), and an additional six were also excluded as the full text was inaccessible (1.08%, reason 3). Furthermore, 102 records were excluded, as they were not academic periodical journal articles (18.28%, reason 4). From the remaining 442 records, an additional four were excluded, as they did not have any abstract available (0.72%, reason 5). Furthermore, a screening based on the presence or absence of keywords [supplementary file] in the full-text was performed, and from this we excluded 123 records that did not include any of our defined keywords (27.83%, reason 6). After the previous screening step, we performed a second filter on the 315 records which were left, based on the same set of keywords as previously but performed on the title and abstract only. A total of 170 (30.05%, reason 7) additional records were then excluded, obtaining a total of 145 records for the final list of included publications from our automatic screening. With the 285 unique records from the manual query, we performed a preliminary exclusion and removed 78 records (27.37%, reason 8). With the remaining 207 articles, we performed our title-based screening and excluded 50 records (17.54%, reason 9), as well as our abstract-based screening, excluding an additional 11 records (3.86%, reason 10). The final content-based screening, based on the full-text of the publications, led to the exclusion of 17 records (5.96%, reason 11). After these screening steps we were left with 129 records. We then removed any articles that were not published in academic periodical journals, leading to the exclusion of 37 records (12.98%, reason 4). We thus obtained the final list of included publications from our manual screening, containing a total of 92 items.
@@
In total we obtained 196 unique items when both results from manual and automatic screening were merged. 42 articles were included through both screening processes, while 51 articles were unique to the manual protocol, and 103 were unique to the automatic one. Additional records were excluded a posteriori during the reviewing of the articles for a wide variety of reasons [Fig. 1, reasons 12, 13, 14 and 15], though this was mostly relevant for those articles that were screened automatically, since the automatic screening script was not exact. In total, 61 records (31.12% of 196) were excluded. The articles that were removed to avoid reviewing articles that the authors contributed to [Fig. 1, reason 14] were McPherron et al. [2022] and Orellana Figueroa et al. [2021]. The final list of articles reviewed contained a total of 135 records.
@@ 3.2 Metadata analysis
Through a chronological analysis of our reviewed articles [Fig. 3], we could distinguish a visible increase in recent years, with over 80% articles reviewed published after 2018. From our quick search for publications between January 2023 and September 2024, which laid outside some of inclusion criteria, and thus we did not review, we obtained a total of 278 unique records. This number is slightly higher (by 12%) than the number of records from 2021 to 2022 from the results of our automatic protocol search (n = 248). In comparison, the number of articles found that were published in 2021 and 2022 found from our automatic protocol search showed an increase of 82% compared to those published in 2019 and 2020 (n = 136). A total of 76 records from the 278 obtained from our 2023 and 2024 search passed the automatic screening protocol. Although publications from the last trimester of 2024 were not included, our results suggest a stabilisation in terms of publication numbers in contrast to the large increase after 2020.
@@
Focusing now only on the 135 publications reviewed, we observed that 10 journals had published nearly half of them (67 records, 49.63%), from a total of 68 different journals observed. Out of these, 6 journals had published nearly a quarter of all articles reviewed [23.70%, Table 3]. The journal Remote Sensing had the most publications, followed by the Journal of Archaeological Science. However, if we use the journal’s impact factor, as well as the journal-wide h-index, multiplying them by the number of publications, we obtain two rudimentary research impact metrics across all the publications reviewed. From this analysis, PLOS One obtained the highest combined score for the h-index-based metric and second for the impact factor–based metric (IF). An analysis of the publication policies revealed that a total of 95 articles (70.37%) were published with an open access modality (including gold, silver, or bronze open access) or were otherwise freely accessible. Some bias towards larger journals or more open access articles may have been occurred due to our search strategy (which for example, did not make use of Scopus) or search queries [see boxes 1 and 2]. Additional bias could have come from a lower interest from these journals in publishing machine learning applications compared to more well-established approaches.
@@
The geographical distribution of the main institutions of the first authors from our set of reviewed articles shows that European and Anglosphere countries, sometimes referred to as the “Global North”, are over-represented [Fig. 4], already observed by Davis [2020a]. Although one could also draw a line between the northern and southern hemispheres to divide the geographical distribution of the analysed publications, this division is not as clear, as countries such as New Zealand and Australia, as well as to some extent Argentina, which is part of the “Global South”, challenge this notion.
@@ 3.3.1 Statistical overview
Among the different subfields of archaeology present in our review, surveying (and site prospection), conservation and cataloguing, as well as classification and typology, are the most represented in our dataset and together account for 49% of all study cases [Fig. 5A]. Landscape archaeology and geoarchaeology are also both well represented with 20 and 17 attributed study cases respectively. The oldest applications (1997 - 2016) of machine learning in archaeology are linked with the subfields of classification and typology as well as conservation and cataloguing. In recent years, there has been a general increase in all topics with no clear trend visible, except for the subfields of surveying as well as conservation and cataloguing, which have considerably increased since 2021 [Fig. 5A]. Only the subfields of zooarchaeology, archaeobotany, and archaeological excavation remain nearly constant through time.
@@
Examining the different architecture of machine learning methods used, we observed that artificial neural networks (ANNs) and ensemble learning represent 62% of all study cases, with 110 and 70 occurrences respectivly [Figs. 2 and 5B]. Linear classifiers, decision trees, rule induction, and nearest neighbour classifiers were also well-represented, accounting for 25% of total applications. An increase of the use of ANNs in our list of reviewed articles from 2021 onwards could be observed, with the number of applications multiplying four-fold compared to the number of applications in the years 2019 and 2020 [Fig. 5B]. At a more granular scale, we found a total of 70 machine learning models [supplementary file]. In addition, we found a particularly high diversity in the unsupervised learning and clustering category of model families, with a unique model for each of the twelve recorded applications for the category [supplementary file]. Despite the dominance of ANNs, random forest was the most common individual model (n = 54). We found that across all articles reviewed, the mean number of models used per study case was 2.12, with a high disparity across different publications, with some applying only one model, while articles such as Bataille et al. [2018] and Courtenay et al. [2019] tested six different models at once. Classification was the most common form of data evaluation, applied in more than 75% of study cases (n = 112). On the other hand, only 19% (n = 29) were regression models, and clustering was only represented with six uses [6%, Figs. 2, 7, 8 and 9].
It was more difficult to observe patterns for the other features we collected, as the attribution to one or several categories for an article was more subjective to the person reviewing. Two of the a posteriori study tasks are highly prominent: automatic structure detection and artefact classification. These two tasks account for 45% (n = 67) of all study case tasks [Fig. 2]. Tasks such as taphonomic classification, archaeological predictive models, and architectural element classification are well represented with 5 to 13 study cases each, whilst nine task categories (e.g. sourcing, species classification, movement recognition) were only relevant for around 5 study cases.
@@
Finally, according to the assessment of article authors or (if not discussed) the results metrics, we could observe that most applications (n = 79) had reported “successful” outcomes [Fig. 6B], with an additional 40 mixed results (27.2%). Only 15 applications (10.2%) were reported to have been “unsuccessful”, whilst we noted 9 (6.1%) study cases whose application had important methodological issues, and classified them as such even if the authors reported a successful outcome. Finally, four (2.7%) study cases did not have a defined outcome. These results, however, could be affected by the well-reported issue of publication bias against the reporting of negative results (Song et al., 2013), as well as the related file-drawer-effect (Rosenthal, 1979; Sparks, 2009), which can strongly affect the perception and production of scientific results (Song et al., 2010).
Figure 5: Fig. 5 (A) Number of articles from each archaeological subfield between 1997 and 2023. (B) Number of articles from each architecture class between 1997 and 2023. Empty bar charts represent the number 1. Figure generated with R 4.2.2 (code available in supplementary material 3).

### v5

@@ 1 Introduction
Unlike traditional statistical models, like linear models such as ordinary least squares or general linear models that find solutions with formal equations, machine learning methods seek to use the data itself to create a model able to accurately evaluate new, unseen data. Machine learning is defined by Alpaydin as “programming computers to optimize a performance criterion using example data or past experience(Alpaydin, 2014, p.3). Machine learning models require a training phase, where the computer learns from the data to improve its model and thus its predictions. Once trained, machine learning models often yield faster, more accurate results than traditional methods, all while lowering the costs [e.g. Kochkov et al. 2024]. This benefit can make machine learning an appealing solution to certain scientific questions (Verhagen, 2007; Hansen and Nebel, 2020; Calder et al., 2022; Brandsen and Koole, 2022; Orellana Figueroa et al., ), Orellana Figueroa et al., in press. However, this comes at the cost of both the models being difficult to interpret (Carvalho et al., 2019), and the training process generally requiring large amounts of (high quality) data, time, and compute power; in a way, front-loading the time and costs to the training process to produce fast and accurate predictions once trained (Sevilla et al., 2022).
@@
The frequency, extent, and success of machine learning applications across archaeology remain largely unknown and their implications and ethical considerations are under debate (Davis, 2020b; Bickler, 2021; Tenzer et al., 2024; Gattiglia, 2025; Neri and Dadà, 2025). Previous reviews of machine learning applications in archaeology mainly focused on remote sensing (Jamil et al., 2022; Argyrou and Agapiou, 2022; Bellat et al., 2025), text analysis (Gonzalez-Perez et al., 2023), classification of ceramics (Ling et al., 2024), wear-use analysis (Eleftheriadou et al., 2025) and other artefacts (Naso and Sciuto, 2025), site-conservation problematics (Casillo et al., 2025), or only on a small set of study cases (Mantovan and Nanni, 2020). On the other hand, while Cacciari and Pocobelli [2022] have assessed machine learning applications across a wide range of fields of archaeological research, they mainly provide qualitative observations rather than a quantitative analysis as presented here, as is also the case for the overview published by Calder et al. [2022]. Palacios [2023] provided an innovative introduction on the status of machine learning in archaeology, although mainly focused on Bayesian approaches, and with a limited selection of reviewed papers mainly on archaeological predictive models. A recent comparison of machine learning methods to PCA and LDA for species identity from faunal material find that methods like RF and LDA outperform PCA (Cole et al., 2022).
@@ 2.1 Search strategy
The search strategy comprised two different and independent protocols: one performed for the automatic screening protocol (considering a fully automated and scripted workflow), consisting of multiple searches, and one performed for the manual screening protocol (mixing automated and human selection steps), consisting of a single search [Table˜1, Fig.˜1]. The searches for the automatic protocol were performed by M.B and consisted of twelve queries each containing a combination of keywords [Box 1]. These keywords, while not covering the entire scope of possible archaeological study cases using machine learning (such as those not using the exact term of "machine learning"), already grasp a wide overview of machine learning applications in the field of archaeology. It is however important to note that this might have biased the search in favour of articles that use the keywords of “machine learning”. On the other hand, an earlier search using the manual protocol involved multiple queries using the 10 machine learning method families we used for our data analysis [see supplementary file], but these returned a large number of duplicate results, most of which were already captured by simply using "machine learning". The searches were performed in five online portals. All searches were conducted in English, except for the German National Library, where the search was conducted in German with only six of the queries [Box 1]. These searches covered all records published before 1 January 2023. In total, 1460 records were retrieved from the automated search protocol, of which 558 were unique.
@@ 2.2 Screening
Following the automatic and manual protocol searches, two independent screening procedures were performed on the retrieved records [Fig.˜1]. For the automatic protocol, after excluding publications not written in English and records not published in scientific journals, the screening was performed with a script written in R 4.4.1 (R Core team, 2024) with the corpus and tm packages (Huang et al., 2021; Feinerer and Hornik, 2023) divided in two filters. The first screening step retrieved all documents whose full text contained one of the archaeology or machine learning keywords we defined for this filtering step [supplementary file]. The second screening step was performed in the same manner with the same keywords, but using only the abstract and title of the results of the first screening step that were found to contain the keywords in the full text. A total of 135 records were included from the automatic search and screening protocol.
The manual screening protocol consisted of additional steps. First, a preliminary exclusion was performed, which checked whether the publications were inaccessible, whether they were preprints that were not yet in print, and whether the record was for an entire book, amongst other criteria [Fig.˜1,supplementary file]. Further screening was then performed based on the exclusion and inclusion criteria [supplementary file], firstly by examining the publication title alone. If after this, an article could not be securely excluded or included, a second screening step followed, now examining the abstract as well. If this was yet still insufficient for a secure inclusion or exclusion, the full-text was examined. A total of 93 records were included from the manual protocol search and screening protocol.
Both screening strategies used the same inclusion and exclusion criteria [Fig.˜1,supplementary file], though the automatic screening script, by necessity, used only a simplified version that only involved the use of keywords, while the manual protocol search was done through the human-based analysis of text.
@@ 2.3 Data extraction
We systematically divided each of the included articles into one or more study cases if different applications (e.g. using different data or seeking different goals) were attested in the publication. Therefore, we obtained a total number of study cases (n = 147) superior to the sum of all included articles (n = 135). From all these study cases, we systematically recorded a total of nine features composed of different categories [Table˜2]. These features were based on the authors’ ad hoc evaluation of possible features of interest in the evaluation of the study cases. No inter-rater reliability was calculated for the selection of these features. However, only four of these were needed to gain a clear overview of the practice of machine learning in archaeology: the architecture of the machine learning model used as well as its evaluation process (see below), the archaeological subfields where these methods were applied, and the type of task it was applied to [Fig.˜2]. Though important, the other groups of features did not allow us to identify common patterns in the application of machine learning in archaeology. Thus, we included them only sporadically in our discussion. Furthermore, we extracted metadata related to the publication for our quantitative analysis. These metadata included the year of publication, the list of authors, the country of affiliation of the first author, the name of the journal, and whether the article was published through an open access modality.
Based on the model architecture, we divided machine learning methods into nine broad categories based on the family of algorithms they belonged to [Eleftheriadou et al. 2025, Fig.˜2, supplementary file]. We granularly recorded the different types of models used, which were then classified into the relevant architecture category, but we also recorded the frequency of their use. Each model was classified into only one architecture category, even in cases where multiple categories may have been warranted. For example, random forest is an ensemble of decision trees, fitting into both “ensemble learning” and “decision trees”, but was classified into “ensemble learning” only, as that was the more relevant aspect of the algorithm.
Furthermore, according to the goals of the model application, we grouped the evaluation into three broad categories: classification, regression, and clustering [Fig.˜2; Alpaydin 2014, pp. 5-13].
@@
Due to the wide range of subjects, we divided the field of archaeology into fifteen subfields [Fig.˜2] based on Kelly and Thomas [2017]. The assignment of a study case to a specific subfield was based on the authors’ own evaluation [supplementary file]. No inter-rater reliability calculation was performed for the assignment. A single publication could be classified into several subfields based on their research questions or objectives, and therefore, the sum total entries in the subfield feature does not equal the total number of study cases included in the review.
Finally, we classified every article’s application into nineteen a posteriori categories based on the broader task (e.g. automatic structure detection) for which the machine learning methods were applied [Fig.˜2]. These task categories were ad hoc and based on the authors’ evaluation of the corpus. Initially, we sought to define the tasks for which machine learning was applied for each publication from a very granular level, then aggregating the more granular task categories into broader ones a total of five times until we arrived at the nineteen categories reported here [supplementary file]. No inter-rater reliability calculation was performed for the creation of these tasks categories. These task categories were one of the most important characteristics to analyse, as these classes helped us understand not only the goals of the authors, but also the possible models that were more well-suited for the desired outcomes, based especially on previous applications.
@@ 3.1 Screening results
Since the two screening protocols were performed independently, we will report them as such, even if the records obtained after both screenings were merged into a single set of publications [Fig.˜1]. Out of the 558 unique records obtained from the automatic protocol, four could not be screened with our script (0.72%, reason 1), another four were excluded, as they were not written in English (0.72%, reason 2), and an additional six were also excluded as the full text was inaccessible (1.08%, reason 3). Furthermore, 102 records were excluded, as they were not academic periodical journal articles (18.28%, reason 4). From the remaining 442 records, an additional four were excluded, as they did not have any abstract available (0.72%, reason 5). Furthermore, a screening based on the presence or absence of keywords [supplementary file] in the full-text was performed, and from this we excluded 123 records that did not include any of our defined keywords (27.83%, reason 6). After the previous screening step, we performed a second filter on the 315 records which were left, based on the same set of keywords as previously but performed on the title and abstract only. A total of 170 (30.05%, reason 7) additional records were then excluded, obtaining a total of 145 records for the final list of included publications from our automatic screening. With the 285 unique records from the manual query, we performed a preliminary exclusion and removed 78 records (27.37%, reason 8). With the remaining 207 articles, we performed our title-based screening and excluded 50 records (17.54%, reason 9), as well as our abstract-based screening, excluding an additional 11 records (3.86%, reason 10). The final content-based screening, based on the full-text of the publications, led to the exclusion of 17 records (5.96%, reason 11). After these screening steps we were left with 129 records. We then removed any articles that were not published in academic periodical journals, leading to the exclusion of 37 records (12.98%, reason 4). We thus obtained the final list of included publications from our manual screening, containing a total of 92 items.
@@
In total we obtained 196 unique items when both results from manual and automatic screening were merged. 42 articles were included through both screening processes, while 51 articles were unique to the manual protocol, and 103 were unique to the automatic one. Additional records were excluded a posteriori during the reviewing of the articles for a wide variety of reasons [Fig.˜1, reasons 12, 13, 14 and 15], though this was mostly relevant for those articles that were screened automatically, since the automatic screening script was not exact. In total, 61 records (31.12% of 196) were excluded. The articles that were removed to avoid reviewing articles that the authors contributed to [Fig.˜1, reason 14] were McPherron et al. [2022] and Orellana Figueroa et al. [2021]. The final list of articles reviewed contained a total of 135 records.
@@ 3.2 Metadata analysis
Through a chronological analysis of the reviewed articles [Fig.˜3], we could distinguish a visible increase in the number of publications in recent years, with over 80% articles reviewed published after 2018. From our quick search for publications between January 2023 and September 2024, which laid outside some of inclusion criteria, and thus we did not review, we obtained a total of 278 unique records. This number is slightly higher (by 12%) than the number of records from 2021 to 2022 from the results of our automatic protocol search (n = 248). In comparison, the number of articles found that were published in 2021 and 2022 found from the automatic protocol search showed an increase of 82% compared to those published in 2019 and 2020 (n = 136). A total of 76 records from the 278 obtained from our 2023 and 2024 search passed the automatic screening protocol. Although publications from the last trimester of 2024 were not included, our results suggest a stabilisation in terms of publication numbers in contrast to the large increase after 2020.
@@
Focusing now only on the 135 publications reviewed, we observed that 10 journals had published nearly half of them (67 records, 49.63%), from a total of 68 different journals observed. Out of these, 6 journals had published nearly a quarter of all articles reviewed [23.70%, Table˜3]. The journal Remote Sensing had the most publications, followed by the Journal of Archaeological Science. However, if we use the journal’s impact factor, as well as the journal-wide h-index, multiplying them by the number of publications, we obtain two rudimentary research impact metrics across all the publications reviewed. From this analysis, PLOS One obtained the highest combined score for the h-index-based metric and second for the impact factor–based metric (IF). An analysis of the publication policies revealed that a total of 95 articles (70.37%) were published with an open access modality (including gold, silver, or bronze) or were otherwise freely accessible. Some bias towards larger journals or more open access articles may have been occurred due to our search strategy (which for example, did not make use of Scopus) or search queries [see boxes 1 and 2]. Additional bias could have come from a lower interest from these journals in publishing machine learning applications compared to more well-established approaches.
@@
The geographical distribution of the main institutions of the first authors from our set of reviewed articles shows that European and Anglosphere countries, sometimes referred to as the “Global North”, are over-represented [Fig.˜4], already observed by Davis [2020a]. Although one could also draw a line between the northern and southern hemispheres to divide the geographical distribution of the analysed publications, this division is not as clear, as countries such as New Zealand and Australia, as well as to some extent Argentina, which is part of the “Global South”, challenge this notion.
@@ 3.3.1 Statistical overview
Among the different subfields of archaeology present in our review, surveying (and site prospection), conservation and cataloguing, as well as classification and typology, are the most represented in our dataset and together account for 49% of all study cases [Fig. 5A]. Landscape archaeology and geoarchaeology are also well represented with 20 and 17 attributed study cases respectively. The oldest applications (1997 - 2016) of machine learning in archaeology are linked with the subfields of classification and typology as well as conservation and cataloguing. In recent years, there has been a general increase in all topics with no clear trend visible, except for the subfields of surveying as well as conservation and cataloguing, which have considerably increased since 2021 [Fig. 5A]. Only the subfields of zooarchaeology, archaeobotany, and archaeological excavation remain nearly constant through time.
@@
Examining the different architecture of machine learning methods used, we observed that artificial neural networks (ANNs) and ensemble learning represent 62% of all study cases, with 110 and 70 occurrences respectivly [Figs. 2 and 5B]. Linear classifiers, decision trees, rule induction, and nearest neighbour classifiers were also well-represented, accounting for 25% of total applications. An increase of the use of ANNs in our list of reviewed articles from 2021 onwards could be observed, with the number of applications multiplying four-fold compared to the number of applications in the years 2019 and 2020 [Fig. 5B]. At a more granular scale, we found a total of 70 machine learning models [Fig. 6 and Annexe 1]. In addition, we found a particularly high diversity in the unsupervised learning and clustering category of model families, with a unique model for each of the twelve recorded applications for the category [Fig. 6 and Annexe 1]. Despite the dominance of ANNs [Fig. 6], random forest was the most common individual model (n = 54). We found that across all articles reviewed, the mean number of models used per study case was 2.12, with a high disparity across different publications, with some applying only one model, while articles such as Bataille et al. [2018] and Courtenay et al. [2019] tested six different models at once. Classification was the most common form of data evaluation, applied in more than 75% of study cases (n = 112). On the other hand, only 19% (n = 29) were regression models, and clustering was only represented with six uses [6%, Figs.˜2, 8, 9 and 10].
It was more difficult to observe patterns for the other features we collected, as the attribution to one or several categories for an article was more subjective to the person reviewing. Two of the a posteriori study tasks are highly prominent: automatic structure detection and artefact classification. These two tasks account for 45% (n = 67) of all study case tasks [Fig.˜2]. Tasks such as taphonomic classification, archaeological predictive models, and architectural element classification are well represented with 5 to 13 study cases each, whilst nine task categories (e.g. sourcing, species classification, movement recognition) were only relevant for around 5 study cases.
@@
Finally, according to the assessment of article authors or (if not discussed) the results metrics, we could observe that most applications (n = 79) had reported “successful” outcomes [Fig. 7B], with an additional 40 mixed results (27.2%). Only 15 applications (10.2%) were reported to have been “unsuccessful”, whilst we noted 9 (6.1%) study cases whose application had important methodological issues, and classified them as such even if the authors reported a successful outcome. Finally, four (2.7%) study cases did not have a defined outcome. These results, however, could be affected by the well-reported issue of publication bias against the reporting of negative results (Song et al., 2013), as well as the related file-drawer effect (Rosenthal, 1979; Sparks, 2009), which can strongly affect the perception and production of scientific results (Song et al., 2010).
Figure 5: (A) Number of articles from each archaeological subfield between 1997 and 2023. (B) Number of articles from each architecture class between 1997 and 2023. Empty bar charts represent the number 1. Figure generated with R 4.2.2 (code available in supplementary material 3).